OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x - 1 < -12 or x + 12 > 10 Answer x < -11 or x > -2 A number line with an open circle on -11, shading to the left, and an open circle on -2, shading to the right. x > -2 or x < -11 A number line with an open circle on -11, an open circle on -2, and shading in between.

jimthompson5910 (jim_thompson5910):

x - 1 < -12 or x + 12 > 10 x < -12+1 or x > 10-12 x < -11 or x > -2 So choice A is out the question ------------------------------------------------------- This graphs to "A number line with an open circle on -11, shading to the left, and an open circle on -2, shading to the right.", so choice B is out ------------------------------------------------------- choice C is just a flip of choice A, so thats out too So the answer is choice D by process of elimination

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 3 < -11 or x + 5 14 Answer A number line with a closed circle on -8, an open circle on 9, and shading in between. A number line with a closed circle on -8, shading to the left, and an open circle on 9, shading to the right. A number line with an open circle on -8, a closed circle on 9, and shading in between. A number line with an open circle on -8, shading to the left, and a closed circle on 9, shading to the right.

jimthompson5910 (jim_thompson5910):

missing a sign...please give me a sign lol

jimthompson5910 (jim_thompson5910):

> or <

OpenStudy (anonymous):

mhm >=

jimthompson5910 (jim_thompson5910):

k thx

OpenStudy (anonymous):

lol sorry agian..

jimthompson5910 (jim_thompson5910):

x - 3 < -11 or x + 5 >= 14 x < -11+3 or x >= 14-5 x < -8 or x >= 9 So we have an open circle at -8 and we shade to the left then we have a closed circle at 9 and we shade to the right so choice D is the answer

jimthompson5910 (jim_thompson5910):

lol it's ok

OpenStudy (anonymous):

thanks jim

jimthompson5910 (jim_thompson5910):

np